OpenStudy (anonymous):

A line has a vertical change per unit of horizontal change of three-halves and passes through point two comma two. Which is another point on the line? A) 5,4 B) 4,6 C) 4,5 D) 6,4 Can You Explain this to me? Thanks!

OpenStudy (fibonaccichick666):

so are you familiar with point slope form of a line?

OpenStudy (anonymous):

Yes but I have trouble with it, but maybe when you explain ill be more familiar.

OpenStudy (fibonaccichick666):

ok, so can you write the formula for me please?

OpenStudy (anonymous):

(x1, y1) and (x 2, y 2) is that right?

OpenStudy (fibonaccichick666):

not quite, but I appreciate you trying :) so the formula is: \[(y-y_0)=m(x-x_0)\]

OpenStudy (anonymous):

your welcome and ok

OpenStudy (fibonaccichick666):

\(y_0~and~x_0\) are the coordinates of the point you are given in this case

OpenStudy (fibonaccichick666):

Do you follow?

OpenStudy (jdoe0001):

\(\bf slope = \cfrac{rise}{run} \implies \cfrac{\text{vertical change}}{\text{horizontal change}}\)

OpenStudy (anonymous):

I just followed you. and ok.

OpenStudy (anonymous):

Thanks for the formula thats easier to understand.

OpenStudy (fibonaccichick666):

So do you know what the m represents?

OpenStudy (anonymous):

The m represents..three-halves ?

OpenStudy (fibonaccichick666):

That is correct in this case, but can you tell me in general what it represents?

OpenStudy (anonymous):

Oh sorry its the horizontal change, correct?

OpenStudy (fibonaccichick666):

not quite, think about the name of the formula

OpenStudy (anonymous):

slope?

OpenStudy (fibonaccichick666):

That is correct.

OpenStudy (anonymous):

Sorry about that I struggle in summer school math

OpenStudy (fibonaccichick666):

So as jdoe supplemented, slope can be described in a lot of ways.

OpenStudy (fibonaccichick666):

it's ok, it happens all of the time

OpenStudy (fibonaccichick666):

So can you guess what our formula will be for the line?

OpenStudy (anonymous):

Yes I saw (: So do I do this.. Slope = 3 over 2 ?

OpenStudy (fibonaccichick666):

yep. :) so can you put that into our formula?

OpenStudy (anonymous):

what do you mean by that?

OpenStudy (fibonaccichick666):

remember the point-slope formula I listed above? That can be used to give the equation of the line. This is something that we need to solve the problem

OpenStudy (fibonaccichick666):

So can you fill in the missing information?

OpenStudy (anonymous):

(3 - 2) =m ?

OpenStudy (fibonaccichick666):

not quite. so let's take a step back, the slope =m=?

OpenStudy (anonymous):

2,2 ?

OpenStudy (fibonaccichick666):

No, what did you say before was the slope?

OpenStudy (anonymous):

3 over 2

OpenStudy (fibonaccichick666):

so m=?

OpenStudy (anonymous):

m = 3 over 2

OpenStudy (fibonaccichick666):

correct so what is \((x_0,y_0)\)? also called any point on the line?

OpenStudy (anonymous):

the point on the line is... (2,2) ?

OpenStudy (fibonaccichick666):

yes! that is correct. now, can you plug the vales into the eq? \[ (y-y_0)=m(x-x_0)\]

OpenStudy (anonymous):

OH! now I get it ok so (2 - 3) =m(2 -2) ?

OpenStudy (fibonaccichick666):

no, no, not quite remember m=?

OpenStudy (anonymous):

oh the other way around like.. (3 -2) =m(2-2) ?

OpenStudy (fibonaccichick666):

no, just answer this. m=?

OpenStudy (anonymous):

3 over 2

OpenStudy (fibonaccichick666):

\(x_0=?\)

OpenStudy (anonymous):

2

OpenStudy (fibonaccichick666):

\(y_0=?\)

OpenStudy (anonymous):

2

OpenStudy (fibonaccichick666):

now exchange the values you just listed for their counterparts in the eq. \[(y-y_0)=m(x-x_0)\]

OpenStudy (anonymous):

(y - 2) =m(x -2) but the net part I dont know yet

OpenStudy (fibonaccichick666):

net part?

OpenStudy (anonymous):

next*

OpenStudy (fibonaccichick666):

so you have not inserted one value, m can you please replace this as well

OpenStudy (anonymous):

I do not know what y & x are yet

OpenStudy (fibonaccichick666):

You don't need those

OpenStudy (fibonaccichick666):

you already know m you just put it in

OpenStudy (anonymous):

Ok so I did the rest right.

OpenStudy (fibonaccichick666):

yes

OpenStudy (anonymous):

Now what?

OpenStudy (fibonaccichick666):

all you do is plug in the corresponding values (x,y) do not have any values. Now write the eq with the substitution for m.

OpenStudy (fibonaccichick666):

so x&y stay x&y

OpenStudy (anonymous):

I dont understand

OpenStudy (fibonaccichick666):

what don't you understand?

OpenStudy (anonymous):

what do I do next? I already did the expression with the (2,2) in and I left Y and X blank so now what..?

OpenStudy (fibonaccichick666):

You literally just put \(\frac{3}{2}\) where the m is in the eq. that's it.

OpenStudy (fibonaccichick666):

you have not replaced it yet

OpenStudy (fibonaccichick666):

that is the eq of the line

OpenStudy (fibonaccichick666):

http://www.purplemath.com/modules/strtlneq.htm for reference

OpenStudy (anonymous):

so.. (3 - 2) =m(2-2) ?

OpenStudy (fibonaccichick666):

NO! what is m=?

OpenStudy (fibonaccichick666):

the m in the eq that you keep typing as an m

OpenStudy (fibonaccichick666):

I said x and y stay there don't rewrite them

OpenStudy (anonymous):

ok then (y - 2) =m(x - 2)

OpenStudy (fibonaccichick666):

replace the m with it's value

OpenStudy (fibonaccichick666):

do not write an m

OpenStudy (anonymous):

so just write 3 and 2 where the 2's are and dont write m

OpenStudy (fibonaccichick666):

no

OpenStudy (fibonaccichick666):

what does m=?

OpenStudy (anonymous):

three-halves

OpenStudy (fibonaccichick666):

so instead of writing m write that.

OpenStudy (anonymous):

ok (y - 2) =3 over 2 (x - 2)

OpenStudy (fibonaccichick666):

yes!

OpenStudy (anonymous):

ok now

OpenStudy (fibonaccichick666):

ok now. just see which points satisfy this eq.

OpenStudy (anonymous):

6,4?

OpenStudy (fibonaccichick666):

so if x=4 what does y=?

OpenStudy (anonymous):

6

OpenStudy (fibonaccichick666):

no,

OpenStudy (fibonaccichick666):

so we get the eq. \(y-2=\frac{3}{2} (4-2)\)

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

just solve

OpenStudy (fibonaccichick666):

what is 4-2=?

OpenStudy (anonymous):

2

OpenStudy (fibonaccichick666):

now what is \(\frac{3}{2}(2)=?\)

OpenStudy (anonymous):

3

OpenStudy (fibonaccichick666):

right. so we have y-2=3

OpenStudy (fibonaccichick666):

so y=?

OpenStudy (anonymous):

5 so its 5,2 ?

OpenStudy (fibonaccichick666):

what?

OpenStudy (fibonaccichick666):

where did you get that?